OpenStudy (anonymous):

find integral of pi/(4+4sin2x)

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/6}\pi/(4+4\sin2x)\]

OpenStudy (anonymous):

\[(\pi*\sin(x))/4(\sin(x)+\cos(x))\] = F(x). Now evaluate evaluate F(x) for that interval.

OpenStudy (anonymous):

i got (pi sqrt3)/6